# Nonasymptotic Analysis of Stochastic Gradient Descent with the Richardson–Romberg Extrapolation

### v2

@@ 1 Introduction
where $\{\gamma_{k}\}_{k\in\mathbb{N}}$ are step sizes, either diminishing or constant, and $\{\xi_{k}\}_{k\in\mathbb{N}}$ is an i.i.d.  sequence with distribution $\mathbb{P}_{\xi}$. The algorithm (2) can be viewed as a special instance of the Robbins-Monro procedure (Robbins & Monro, 1951). While the SGD algorithm remains one of the core algorithms in statistical inference, its performance can be enhanced by means of additional techniques that use e.g., momentum (Qian, 1999), averaging (Ruppert, 1988; Polyak & Juditsky, 1992), or variance reduction (Defazio et al., 2014; Nguyen et al., 2017). In particular, the celebrated Polyak-Ruppert algorithm proceeds with a trajectory-wise averaging of the estimates
@@
[…]
@@
For $\theta_{1},\ldots,\theta_{k}$ being the iterates of stochastic first-order method, we denote ${\mathcal{F}}_{k}=\sigma(\theta_{0},\theta_{1},\ldots,\theta_{k})$. Let $(\mathsf{Z},\mathsf{d}_{\mathsf{Z}})$ be a complete separable metric space equipped with its Borel $\sigma$-algebra $\mathcal{Z}$. We call a function $c:\mathsf{Z}\times\mathsf{Z}\rightarrow\mathbb{R}_{+}$ a distance-like function, if it is symmetric, lower semi-continuous and $c(x,y)=0$ if and only if $x=y$, and there exists $q\in\mathbb{N}$ such that $(d(x,y)\wedge 1)^{q}\leq c(x,y)$. We denote by $\mathscr{C}(\xi,\xi^{\prime})$ the set of couplings of probability measures $\xi$ and $\xi^{\prime}$, that is, a set of probability measures on $(\mathsf{Z}^{2},\mathcal{Z}^{\otimes 2})$, such that for any $\Pi\in\mathscr{C}(\xi,\xi^{\prime})$ and any $A\in\mathcal{Z}$ it holds $\Pi(\mathsf{Z}\times A)=\xi^{\prime}(A)$ and $\Pi(A\times\mathsf{Z})=\xi(A)$. We define the Wasserstein semimetric associated to the distance-like function $c(\cdot,\cdot)$, as
@@ 2 Literature review
Constant step-size SGD has been widely studied in the literature. Its bias and MSE were studied for strongly convex problems in (Dieuleveut et al., 2020), for both the last iterate and the Polyak-Ruppert averaging. Yu et al. (2021) studied the bias and the asymptotic normality of the last iterate of SGD for non-convex problems under the Polyak-Lojasiewicz condition and its generalizations. Vlatakis-Gkaragkounis et al. (2024) considered constant step-size methods for solving variational inequalities, focusing on characterizing the bias and establishing asymptotic normality. Merad & Gaïffas (2023) studied the convergence of constant step-size SGD iterates in $\mathbf{W}_{{p}}$ distance, $p\geq 1$, and provided concentration bounds for both $\theta_{n}$ and $\bar{\theta}_{n_{0},n}$. Moulines & Bach (2011) derived the root-MSE bound (5) with $\delta=1/6$ for the case of strongly convex functions $f$. Gadat & Panloup (2023) derived an MSE counterpart to (5) of the form
@@
Richardson-Romberg (RR) extrapolation is a technique used to improve the accuracy of numerical approximations (Hildebrand, 1987), such as those from numerical differentiation or integration. It involves using approximations with different step sizes and then extrapolating to reduce the error, typically by removing the leading term in the error expansion. The one-step RR extrapolation was introduced to reduce the discretization error induced by an Euler scheme to simulate stochastic differential equation in Talay & Tubaro (1990), and later generalized for non-smooth functions in Bally & Talay (1996). This technique was extended using multistep discretizations in Pagès (2007). RR extrapolation have been applied to Stochastic Gradient Descent (SGD) methods in Durmus et al. (2016), Merad & Gaïffas (2023) and Huo et al. (2024a), to improve convergence and reduce error in optimization problems, particularly when dealing with noisy or high-variance gradient estimates. Recent papers (Allmeier & Gast, 2024; Zhang & Xie, 2024; Huo et al., 2024a; Kwon et al., 2024) consider applications of RR to different stochastic approximation problems with constant step-size, including $Q$-learning, and single- and two-timescale stochastic approximation.
@@ 3.1 Geometric ergodicity of SGD iterates
We aim to solve the problem (1) using SGD with a constant step size, starting from initial distribution $\nu$. That is, for $k\geq 0$ and a step size $\gamma\geq 0,$ we consider the following recurrent scheme
@@
where $\{\xi_{k}\}_{k\in\mathbb{N}}$ is a sequence satisfying the following condition.
@@
and trace only dependence upon $\operatorname{L}$ in our subsequent bounds. In this paper we focus on the convergence to $\theta^{\star}$ of the Polyak-Ruppert averaging estimator defined for any $n\geq 0$,
@@
Many previous studies instead consider $\bar{\vartheta}_{n}^{(\gamma)}=\frac{1}{n-n_{0}}\sum_{k=n_{0}+1}^{n}\theta_{k}^{(\gamma)}$ rather than $\bar{\theta}_{n}^{(\gamma)}$, where $n\geq n_{0}+1$ and $n_{0}$ denotes a burn-in period. However, when the sample size $n$ is sufficiently large, the choice of the optimal burn-in size $n_{0}$ affects the leading terms in the MSE bound of $\bar{\theta}_{n}^{(\gamma)}-\theta^{\star}$ only by a constant factor. Therefore, we focus on (14), or equivalently, use $2n$ observations and set $n_{0}=n$.
@@
where $\mathsf{B}(\mathbb{R}^{d})$ is a Borel $\sigma$-field of $\mathbb{R}^{d}$. In Dieuleveut et al. (2020) it has been established that, under the stated assumptions, ${\rm Q}_{\gamma}$ admits a unique invariant distribution $\pi_{\gamma}$, if $\gamma$ is small enough. Previous studies, such as Dieuleveut et al. (2020) or Merad & Gaïffas (2023), studied the convergence of the distributions of $\{\theta_{k}^{(\gamma)}\}_{k\in\mathbb{N}}$ to $\pi_{\gamma}$ in the $p$-Wasserstein distance $\mathbf{W}_{{p}}$, $p\geq 1$, associated with the Euclidean distance in $\mathbb{R}^{d}$. Our main results require to switch to the non-standard distance-like function, which is defined under A 1 and A 3($2$) as follows:
$c(\theta,\theta^{\prime})=\|\theta-\theta^{\prime}\|\left(\|\theta-\theta^{*}\|+\|\theta^{\prime}-\theta^{*}\|+\frac{2\sqrt{2}\tau_{2}\sqrt{\gamma}}{\sqrt{\mu}}\right)\,,\qquad\theta,\theta^{\star}\in\mathbb{R}^{d}\,.$ (16)
@@ Discussion.
The property (19) is crucial to obtain tighter (with respect to the step size $\gamma$) error bounds for the Richardson-Romberg estimator, as well as in the Rosenthal inequality for additive functional of $\{\theta_{k}^{(\gamma)}\}_{k\in\mathbb{N}}$ derived in Section 4. Precisely, the additional $\sqrt{\gamma}$ factor obtained in (19) would allow us to obtain sharper bounds on the remainder terms in Theorem 2. Next, we analyze the error $\theta_{\infty}^{(\gamma)}-\theta^{\star}$ where $\theta_{\infty}^{(\gamma)}$ is distributed according to the stationary distribution $\pi_{\gamma}$. To this end, we consider the following condition.
@@
It is important to note that C 1 is not independent from the preceding assumptions A 1 - A 3($p$). In particular, Dieuleveut et al. (2020, Lemma $13$) establishes that, under A 1,A 2, and A 3($p$), the bound (21) holds for $\gamma\in(0,1/(\operatorname{L}\mathsf{C}_{\operatorname{step},p})]$ with some constants $\mathsf{D}_{\operatorname{last},p}$ and $\mathsf{C}_{\operatorname{step},p}$ depending only upon $p$. Unfortunately, it is difficult to obtain precise dependence of $\mathsf{C}_{\operatorname{step},p}$ and $\mathsf{D}_{\operatorname{last},p}$ on $p$, and to obtain (21) with precise numerical constants. Existing studies (Gadat & Panloup, 2023; Merad & Gaïffas, 2023) either use different set of assumptions or do not explicitly characterize their dependence on $p$. That is why we prefer to state C 1($p$) as a separate assumption. In the subsequent bounds we use C 1($p$) together with A 1,A 2, and A 3($p$), tracking the dependence of our bounds upon $\mathsf{C}_{\operatorname{step},p}$ and $\mathsf{D}_{\operatorname{last},p}$. We leave the derivation of C 1($p$) with precise constants $\mathsf{D}_{\operatorname{last},p},\mathsf{C}_{\operatorname{step},p}$ as an interesting direction for future research.
Under assumption C 1, we control the fluctuations of $\theta_{k}^{(\gamma)}$ around $\theta^{\star}$. However, unless $f$ is quadratic, it is known that $\int_{\mathbb{R}^{d}}\theta\pi_{\gamma}(\mathrm{d}\theta)\neq\theta^{\star}$. In the next proposition, we quantify this bias under under weaker assumptions than those in Dieuleveut et al. (2020, Theorem 4).
[…]
@@
Here $B_{1}\in\mathbb{R}^{d}$ and $B_{2}\in\mathbb{R}^{d\times d}$ satisfy $\|B_{1}\|\leq\frac{L}{\mu}\mathsf{C}_{1}+\frac{L\mathsf{D}_{\operatorname{last},6}^{3/2}\tau_{6}^{3}}{2\mu^{5/2}}$, $\|B_{2}\|\leq\mathsf{C}_{1}$, where $\mathsf{C}_{1}$ defined in (51) is a constant independent of $\gamma$. Moreover, for any initial distribution $\nu$ on $\mathbb{R}^{d}$, it holds that
@@
In this section, we analyze the finite-sample properties of the estimator $\bar{\theta}_{n}^{(\gamma)}$ from (14). The analysis is based on techniques previously used in Moulines & Bach (2011), as well as in the analysis of the Polyak-Ruppert averaged LSA (Linear Stochastic Approximation) algorithms, see Mou et al. (2020); Durmus et al. (2024). Below we derive the key representation for the error $\bar{\theta}_{n}^{(\gamma)}-\theta^{\star}$, following Moulines & Bach (2011). Define the $k$-th step noise level at the point $\theta\in\mathbb{R}^{d}$ by:
@@
and ${\varepsilon_{k+1}}(\theta_{k}^{(\gamma)})$ is a martingale-difference sequence w.r.t. $({\mathcal{F}}_{k})_{k\in\mathbb{N}}$. Then the recurrence (11) takes form
@@
$\operatorname{H}^{\star}(\theta_{k}^{(\gamma)}-\theta^{\star})=\frac{\theta_{k}^{(\gamma)}-\theta_{k+1}^{(\gamma)}}{\gamma}-{\varepsilon_{k+1}}(\theta_{k}^{(\gamma)})-\eta(\theta_{k}^{(\gamma)})\,.$ (29)
@@
$\operatorname{H}^{\star}(\bar{\theta}_{n}^{(\gamma)}-\theta^{\star})=\frac{\theta_{n+1}^{(\gamma)}-\theta^{\star}}{\gamma n}-\frac{\theta_{2n+1}^{(\gamma)}-\theta^{\star}}{\gamma n}-\frac{1}{n}\sum_{k=n+1}^{2n}{\varepsilon_{k+1}}(\theta_{k}^{\gamma})-\frac{1}{n}\sum_{k=n+1}^{2n}\eta(\theta_{k}^{\gamma})\,.$ (30)
Equation (30) is the key ingredient in the proof of the next result where the variance of noise ${\varepsilon_{k}}(\theta^{\star})$ measured at the optimal point $\theta^{\star}$ naturally appears, that is,
@@
Note that $\Sigma_{\varepsilon}^{\star}$ does not depend on the step size $\gamma$ of (11) and is related to the ”optimal” covariance matrix of the Polyak-Ruppert averaged iterates $\bar{\theta}_{n}^{(\gamma)}$, see Fort (2015). In our first result below, we establish root MSE bound on the error of the Polyak-Ruppert averaged estimator (14).
@@
Assume A 1, A 2, A 3($6$), and C 1($6$). Then for any $\gamma\in(0,1/(\operatorname{L}\mathsf{C}_{\operatorname{step},6})]$, $n\in\mathbb{N}$, and initial distribution $\nu$ on $\mathbb{R}^{d}$, the sequence of Polyak-Ruppert estimates (14) satisfies
@@
The version of Theorem 1 with explicit constants together with the proof is provided in Appendix B, see Theorem 4. Note that the result of Theorem 1 is valid for arbitrary $\gamma\in(0,1/(\operatorname{L}\mathsf{C}_{\operatorname{step},6})]$. At the same time, this bound can be optimized over step size of the form $\gamma=n^{-\beta}$, $\beta\in(0,1)$.
[…]
Under the assumptions of Theorem 1, provided that $n\geq\left(\operatorname{L}\mathsf{C}_{\operatorname{step},6}\right)^{3/2}$, it holds setting $\gamma=n^{-2/3}$ that
@@
Section 3.2 implies that, if $n$ is known in advance and $\gamma=n^{-2/3}$, then $\bar{\theta}_{n}^{(\gamma)}$ satisfies (5) with $\delta=1/6$. A closer inspection of the sum (30) reveals that $\mathsf{E}_{\pi_{\gamma}}[\eta(\theta_{k}^{\gamma})]$ is of order $\gamma$, and we can not expect to provide a better bound for the term $\frac{1}{n}\sum_{k=n+1}^{2n}\eta(\theta_{k}^{\gamma})$ compared to the one coming from Minkowski’s inequality. Thus, this is the bias of the stationary distribution, which does not allows us to improve scaling of the second-order term w.r.t. the sample size $n$.

### v3

@@ 1 Introduction
where $\{\gamma_{k}\}_{k\in\mathbb{N}}$ are step sizes, either diminishing or constant, and $\{\xi_{k}\}_{k\in\mathbb{N}}$ is a sequence of i.i.d. random variables with distribution $\mathbb{P}_{\xi}$. The algorithm (2) can be viewed as a special instance of the Robbins-Monro procedure (Robbins & Monro, 1951). While the SGD algorithm remains one of the core algorithms in statistical inference, its performance can be enhanced by means of additional techniques that use e.g., momentum (Qian, 1999), averaging (Ruppert, 1988; Polyak & Juditsky, 1992), or variance reduction (Defazio et al., 2014; Nguyen et al., 2017). In particular, the celebrated Polyak-Ruppert algorithm proceeds with a trajectory-wise averaging of the estimates
@@
$\sqrt{n}(\bar{\theta}_{n_{0},n}-\theta^{\star})\overset{d}{\to}\mathrm{N}(0,\Sigma_{\infty})\,,\quad n\to\infty$ (4)
where $\overset{d}{\to}$ denotes the convergence in distribution and $\mathrm{N}(0,\Sigma_{\infty})$ denotes the zero-mean Gaussian distribution with covariance matrix $\Sigma_{\infty}$, which is asymptotically optimal, see Fort (2015) for a discussion. On the other hand, quantitative counterparts of (4) rely on the root mean squared error (root-MSE) bounds of the form
@@
For $\theta_{0},\ldots,\theta_{k}$ being the iterates of stochastic first-order method, we denote ${\mathcal{F}}_{k}=\sigma(\theta_{0},\theta_{1},\ldots,\theta_{k})$. Let $(\mathsf{Z},\mathsf{d}_{\mathsf{Z}})$ be a complete separable metric space equipped with its Borel $\sigma$-algebra $\mathcal{Z}$. We call a function $c:\mathsf{Z}\times\mathsf{Z}\rightarrow\mathbb{R}_{+}$ a distance-like function, if it is symmetric, lower semi-continuous and $c(x,y)=0$ if and only if $x=y$, and there exists $q\in\mathbb{N}$ such that $\mathsf{d}_{\mathsf{Z}}^{q}(x,y)\leq c(x,y)$. We denote by $\mathscr{C}(\xi,\xi^{\prime})$ the set of couplings of probability measures $\xi$ and $\xi^{\prime}$, that is, a set of probability measures on $(\mathsf{Z}^{2},\mathcal{Z}^{\otimes 2})$, such that for any $\Pi\in\mathscr{C}(\xi,\xi^{\prime})$ and any $A\in\mathcal{Z}$ it holds $\Pi(\mathsf{Z}\times A)=\xi^{\prime}(A)$ and $\Pi(A\times\mathsf{Z})=\xi(A)$. We define the Wasserstein semimetric associated to the distance-like function $c(\cdot,\cdot)$, as
@@ 2 Literature review
Constant step-size SGD has been widely studied in the literature. Its bias and MSE were studied for strongly convex problems in (Dieuleveut et al., 2020), for both the last iterate and the Polyak-Ruppert averaging. Yu et al. (2021) studied the bias and the asymptotic normality of the last iterate of SGD for non-convex problems under the Polyak-Lojasiewicz condition and its generalizations. Vlatakis-Gkaragkounis et al. (2024) considered constant step size methods for solving variational inequalities, characterizing the bias and establishing asymptotic normality. Merad & Ga¨ıffas (2023) studied the convergence of constant step-size SGD iterates in $\mathbf{W}_{{p}}$ distance, $p\geq 1$, and provided concentration bounds for both $\theta_{n}$ and $\bar{\theta}_{n_{0},n}$. Moulines & Bach (2011) derived the root-MSE bound (5) with $\delta=1/6$ for the case of strongly convex functions $f$. Gadat & Panloup (2023) derived an MSE counterpart to (5) of the form
@@
Richardson-Romberg (RR) extrapolation is a technique used to improve the accuracy of numerical approximations (Hildebrand, 1987), such as those from numerical differentiation or integration. It involves using approximations with different step sizes and then extrapolating to reduce the error, typically by removing the leading term in the error expansion. The one-step RR extrapolation was introduced to reduce the discretization error induced by an Euler scheme to simulate stochastic differential equation in Talay & Tubaro (1990), and later generalized for non-smooth functions in Bally & Talay (1996). This technique was extended using multistep discretizations in Pagès (2007). RR extrapolation have been applied to Stochastic Gradient Descent (SGD) methods in Durmus et al. (2016), Dieuleveut et al. (2020), Merad & Ga¨ıffas (2023) and Huo et al. (2024a), to improve convergence and reduce error in optimization problems, particularly when dealing with noisy or high-variance gradient estimates. Recent papers (Allmeier & Gast, 2024; Zhang & Xie, 2024; Huo et al., 2024a; Kwon et al., 2024) consider applications of RR to different stochastic approximation problems with constant step-size, including $Q$-learning, and single- and two-timescale stochastic approximation.
@@ 3.1 Geometric ergodicity of SGD iterates
We aim to solve the problem (1) using SGD with a constant step size, starting from initial distribution $\nu$. That is, for $k\geq 0$ and a step size $\gamma>0,$ we consider the following recurrent scheme
@@
where $\{\xi_{k}\}_{k\in\mathbb{N}}$ satisfy the following condition:
@@
and trace only dependence upon $\operatorname{L}$ in our subsequent bounds. In this paper we focus on the convergence to $\theta^{\star}$ of the Polyak-Ruppert averaged estimator defined for any $n\geq 0$, by
@@
Many previous studies instead consider $\bar{\theta}_{n_{0},n}=\frac{1}{n-n_{0}}\sum_{k=n_{0}+1}^{n}\theta_{k}^{(\gamma)}$ rather than $\bar{\theta}_{n}^{(\gamma)}$, where $n\geq n_{0}+1$ and $n_{0}$ denotes a burn-in period. However, when the sample size $n$ is sufficiently large, the choice of the optimal burn-in size $n_{0}$ affects the leading terms in the MSE bound of $\bar{\theta}_{n}^{(\gamma)}-\theta^{\star}$ only by a constant factor. Therefore, we focus on (14), or equivalently, use $2n$ observations and set $n_{0}=n$.
@@
where $\mathsf{B}(\mathbb{R}^{d})$ is a Borel $\sigma$-field of $\mathbb{R}^{d}$. In Dieuleveut et al. (2020) it has been established that, under the stated assumptions, ${\rm Q}_{\gamma}$ admits a unique invariant distribution $\pi_{\gamma}$, if $\gamma$ is small enough. Previous studies, such as Dieuleveut et al. (2020) or Merad & Ga¨ıffas (2023), studied the convergence of the distributions of $\{\theta_{k}^{(\gamma)}\}_{k\in\mathbb{N}}$ to $\pi_{\gamma}$ in the $p$-Wasserstein distance $\mathbf{W}_{{p}}$, $p\geq 1$, associated with the Euclidean distance in $\mathbb{R}^{d}$. Our main results require to switch to the non-standard distance-like function, which is defined under A 1 and A 3($2$) as follows:
$c(\theta,\theta^{\prime})=\|\theta-\theta^{\prime}\|\left(\|\theta-\theta^{*}\|+\|\theta^{\prime}-\theta^{*}\|+\frac{2\sqrt{2}\tau_{2}\sqrt{\gamma}}{\sqrt{\mu}}\right)\,,\qquad\theta,\theta^{\prime}\in\mathbb{R}^{d}\,.$ (16)
@@ Discussion.
The property (19) is crucial to obtain tighter error bounds with respect to the step size $\gamma$ for the Richardson-Romberg estimator, as well as to obtain the improved version of the Rosenthal inequality for additive functional of $\{\theta_{k}^{(\gamma)}\}_{k\in\mathbb{N}}$ derived in Section 4. Precisely, the additional $\sqrt{\gamma}$ factor obtained in (19) would allow us to obtain sharper bounds on the remainder terms in Theorem 6. Next, we analyze the error $\theta_{\infty}^{(\gamma)}-\theta^{\star}$ where $\theta_{\infty}^{(\gamma)}$ is distributed according to the stationary distribution $\pi_{\gamma}$. To this end, we consider the following condition.
@@
It is important to note that C 1 is not independent from the preceding assumptions A 1 - A 3($p$). In particular, Dieuleveut et al. (2020, Lemma $13$) establishes that, under A 1,A 2, and A 3($p$), the bound (21) holds for $\gamma\in(0,1/(\operatorname{L}\mathsf{C}_{\operatorname{step},p})]$ with some constants $\mathsf{D}_{\operatorname{last},p}$ and $\mathsf{C}_{\operatorname{step},p}$ depending only upon $p$. Unfortunately, it is difficult to obtain precise dependence of $\mathsf{C}_{\operatorname{step},p}$ and $\mathsf{D}_{\operatorname{last},p}$ on $p$, and to obtain (21) with precise numerical constants. Existing studies (Gadat & Panloup, 2023; Merad & Ga¨ıffas, 2023) either use different set of assumptions or do not explicitly characterize their dependence on $p$. That is why we prefer to state C 1($p$) as a separate assumption. In the subsequent bounds we use C 1($p$) together with A 1,A 2, and A 3($p$), tracking the dependence of our bounds upon $\mathsf{C}_{\operatorname{step},p}$ and $\mathsf{D}_{\operatorname{last},p}$. We leave the derivation of C 1($p$) with precise constants $\mathsf{D}_{\operatorname{last},p},\mathsf{C}_{\operatorname{step},p}$ as an interesting direction for future research.
Under assumption C 1, we control the fluctuations of $\theta_{k}^{(\gamma)}$ around $\theta^{\star}$. However, unless $f$ is quadratic, it is known that $\int_{\mathbb{R}^{d}}\theta\pi_{\gamma}(\mathrm{d}\theta)\neq\theta^{\star}$. In the next proposition, we quantify this bias under weaker assumptions than those in Dieuleveut et al. (2020, Theorem 4).
Proposition 2.
@@
Here $B_{1}\in\mathbb{R}^{d}$ and $B_{2}\in\mathbb{R}^{d\times d}$ satisfy $\|B_{1}\|\leq\mathsf{C}_{b,1}$, $\|B_{2}\|\leq\mathsf{C}_{b,2}$, where $\mathsf{C}_{b,1}$ and $\mathsf{C}_{b,2}$ are constants independent of $\gamma$ and are defined in (LABEL:eq:_def_ConstPRb1,2). Moreover, for any initial distribution $\nu$ on $\mathbb{R}^{d}$, it holds that
@@
In this section, we analyze the finite-sample properties of the estimator $\bar{\theta}_{n}^{(\gamma)}$ from (14). The analysis is based on techniques previously used in Moulines & Bach (2011), as well as in the analysis of the Polyak-Ruppert averaged LSA (Linear Stochastic Approximation) algorithms, see Mou et al. (2020); Durmus et al. (2025). Below we derive the key representation for the error $\bar{\theta}_{n}^{(\gamma)}-\theta^{\star}$, following Moulines & Bach (2011). Define the $k$-th step noise at the point $\theta\in\mathbb{R}^{d}$ by:
@@
and note that ${\varepsilon_{k+1}}(\theta_{k}^{(\gamma)})$ is a martingale-difference sequence w.r.t. $({\mathcal{F}}_{k})_{k\in\mathbb{N}}$. Then the recurrence (11) takes form
@@
$\operatorname{H}^{\star}(\theta_{k}^{(\gamma)}-\theta^{\star})=\gamma^{-1}(\theta_{k}^{(\gamma)}-\theta_{k+1}^{(\gamma)})-{\varepsilon_{k+1}}(\theta_{k}^{(\gamma)})-\eta(\theta_{k}^{(\gamma)})\,.$ (29)
@@
$\operatorname{H}^{\star}(\bar{\theta}_{n}^{(\gamma)}-\theta^{\star})=\frac{\theta_{n+1}^{(\gamma)}-\theta^{\star}}{\gamma n}-\frac{\theta_{2n+1}^{(\gamma)}-\theta^{\star}}{\gamma n}-\frac{1}{n}\sum_{k=n+1}^{2n}{\varepsilon_{k+1}}(\theta_{k}^{(\gamma)})-\frac{1}{n}\sum_{k=n+1}^{2n}\eta(\theta_{k}^{(\gamma)})\,.$ (30)
We further introduce the covariance matrix of ${\varepsilon_{k}}(\theta^{\star})$ measured at the optimal point $\theta^{\star}$, that is,
@@
Note that $\Sigma_{\varepsilon}^{\star}$ does not depend on the step size $\gamma$ and is related to the asymptotically optimal covariance matrix of the Polyak-Ruppert averaged iterates $\bar{\theta}_{n}^{(\gamma)}$, see Fort (2015). Precisely, under assumptions A 1-A 3, the asymptotic covariance matrix $\Sigma_{\infty}$ from (4) is given by
$\Sigma_{\infty}=\{\operatorname{H}^{\star}\}^{-1}\Sigma_{\varepsilon}^{\star}\{\operatorname{H}^{\star}\}^{-1}\,.$ (32)
We state our subsequent results in terms of $\operatorname{H}^{\star}(\bar{\theta}_{n}^{(\gamma)}-\theta^{\star})$ and its Richardson-Romberg adjusted counterpart. One can switch to the corresponding results for $\bar{\theta}_{n}^{(\gamma)}-\theta^{\star}$ at the price of the factor $\|\{\operatorname{H}^{\star}\}^{-1}\|$, which affects the non-leading terms. In our first result below, we establish the root MSE bound on the error of the Polyak-Ruppert averaged estimator (14).
@@
Assume A 1, A 2, A 3($4$), and C 1($4$). Then for any $\gamma\in(0,1/(\operatorname{L}\mathsf{C}_{\operatorname{step},4})]$, $n\in\mathbb{N}$, and initial distribution $\nu$ on $\mathbb{R}^{d}$, the sequence of Polyak-Ruppert estimates (14) satisfies
@@
The version of Theorem 3 with explicit constants together with the proof is provided in Appendix B, see Theorem 15. Note that the result of Theorem 3 is valid for arbitrary $\gamma\in(0,1/(\operatorname{L}\mathsf{C}_{\operatorname{step},4})]$. At the same time, this bound can be optimized over step size of the form $\gamma=n^{-\beta}$, $\beta\in(0,1)$.
Corollary 4.
Under the assumptions of Theorem 3, provided that $n\geq\left(\operatorname{L}\mathsf{C}_{\operatorname{step},4}\right)^{3/2}$, it holds setting $\gamma=n^{-2/3}$ that
@@
Section 3.2 implies that, if $n$ is known in advance and $\gamma=n^{-2/3}$, then $\bar{\theta}_{n}^{(\gamma)}$ satisfies (5) with $\delta=1/6$. A closer inspection of the sum (30) reveals that $\mathsf{E}_{\pi_{\gamma}}[\eta(\theta_{k}^{(\gamma)})]$ is of order $\gamma$, and we can not expect to provide a better bound for the term $\frac{1}{n}\sum_{k=n+1}^{2n}\eta(\theta_{k}^{(\gamma)})$ compared to the one coming from Minkowski’s inequality. Thus, this is the bias of the stationary distribution, which does not allow us to improve scaling of the second-order term w.r.t. the sample size $n$.